PROBLEM 1 i A medication label and accompanying insert contain the following information: Pfizerpen penicillin G potassium for injection twenty million units For intravenous infusion only Atter reconstitution, solution should be refrigerated. Discard unused solution after 7 days. diluent added approx, concentration 75 mL 250,000 units/mL 33 mL 500,000 units/mL 11.5 mL 1,000,000 units/mL 1.What is the expiration date and time if this medication is reconstituted on May 27 at 14152 June sid ad 1415 If you add 75 mL of diluent to the vial, what volume will you draw up for a dose of 2.5 million units? Can this medication be given by intramuscular injection? No, Only inhoenous iafusion only If 11.5 mL of diluent is added to the vial, what volume will you draw up for a dose of 4 million units
Answer #1

1. The solution should be discarded after 7 days of reconstitution.
The date and time of reconstitution is May 27 at 1415
The date and time of expiry is date and time of reconstitution + 7 days
May has 31 days so the expiry date and time is June 3rd 1415

2. 75 ml of diluent yield approximately 250000 units / ml
1 ml = 250000 units
2.5 million = 2500000 units
Formula = desired amount / dose on hand = 2500000 / 250000 = 10
Answer is 10 ml

3. It is mentioned in the label that this injection is only for intravenous infusion.

4. 11.5 ml of diluent yield approximately 1000000 unit / ml
Desired amount = 4 million units I.e. 4000000 units
Formula = desired amount / amount on hand = 4000000 / 1000000 = 4 ml
Answer = 4 ml